OpenStudy (anonymous):

Elliott’s backyard is 4800 square feet. His dad is going to pave 15% of this area to use as a patio. How many square feet will be paved?

OpenStudy (anonymous):

100% - 15% = 85% = 0,85 0,85 * 4800 square feet = 4080 square feet

OpenStudy (mathsolver):

The question asks how many square feet will be paved. So the answer is simply 4800*.15=720

OpenStudy (anonymous):

Oops, my bad!

OpenStudy (anonymous):

thanks for the awnser been on this site long

OpenStudy (anonymous):

what is 26% of 160

OpenStudy (anonymous):

0,26 * 160 = 41,6

OpenStudy (mathsolver):

just multiply 160 by .26 so its 41.6

OpenStudy (anonymous):

what percent of 90 is 18

OpenStudy (anonymous):

18/90 = 1/5 = 0,2 = 20%

OpenStudy (anonymous):

a friend lives in a town 160 miles away ive droven 64 miles how much more do i have to drive

OpenStudy (anonymous):

You have to drive 160 - 64 = 96 miles, that is 96/160 = 3/5 = 0,6 = 60%.

OpenStudy (anonymous):

what is 12.5 percent of 72

OpenStudy (anonymous):

0,125 * 72 = 9

OpenStudy (anonymous):

15% of what number is 24

OpenStudy (anonymous):

24/0,15 = 160.

OpenStudy (anonymous):

35% of 220

OpenStudy (anonymous):

0,35 * 220 = 77 You should've learned by now, come on :D

OpenStudy (anonymous):

be back thursday

OpenStudy (anonymous):

You had a test or something?

OpenStudy (anonymous):

sort of